# A Large and Balanced Corpus for Fine-grained Arabic Readability Assessment

### v1

@@ 6.3 Stage 2 and Ensemble Results
To further improve performance, we experiment with ensemble methods. We define the Average ensemble, where the final prediction is the rounded average of the levels predicted by the six models, and the Most Common ensemble, where the final prediction is the mode of the predicted levels. The results show that the Average ensemble performs better in terms of Acc $\pm$1, Dist, and QWK, indicating that it tends to stay closer to the correct label. However, it struggles with exact accuracy (Acc), as averaging can blur distinctions between classes. On the other hand, the Most Common ensemble achieves higher Acc but can sometimes be misled by an incorrect majority, leading to greater deviation from the correct label.
@@
We also report an Oracle combination, where we assume access to the best possible prediction from the six models for each sample. This serves as an upper bound on model performance. The Oracle results are significantly higher than those of individual models and are comparable to human annotators’ agreement with the unified labels (see section 6.1). This suggests that while individual models are still far from human-level performance, ensembling has the potential to push results closer to human agreement. More oracle combinations are provided in Appendix E.3. We also include more results on the impact of training granularity on readability level prediction in Appendix E.2 Finally, we note that the trends observed in the development set persist in the test set, further validating our findings.
@@ 7 Conclusions and Future Work
This paper presented the Balanced Arabic Readability Evaluation Corpus (BAREC), a large-scale, finely annotated dataset for assessing Arabic text readability across 19 levels. With over 68,000 sentences and 1 million words, it is the largest Arabic corpus for readability assessment, covering diverse genres, topics, and audiences, to our knowledge. High inter-annotator agreement ensures reliable annotations. Through benchmarking various readability assessment techniques, we highlighted both the challenges and opportunities in Arabic readability modeling, demonstrating promising performance across different methods. Looking ahead, we plan to expand the corpus, enhancing its size and diversity to cover additional genres and topics. We also aim to add annotations related to vocabulary leveling and syntactic treebanks to study less-explored genres in syntax. Future work will include analyzing readability differences across genres and topics. Additionally, the tools we have developed will be integrated into a system to help children’s story writers target specific reading levels. The Barec dataset, its annotation guidelines, and benchmark results, will be made publicly available to support future research and educational applications in Arabic readability assessment.
@@ C.1 Resources
A subset of 277 books from Hindawi accross different different genres.  [^4] https://www.hindawi.org/books/categories/
Wikipedia
A subset of 175 Arabic wikipedia articles covering Culture, Figures, Geography, History, Mathematics, Sciences, Society, Philosophy, Religions and Technologies. [^5] https://ar.wikipedia.org/
ChatGPT
To add more children’s materials, we ask Chatgpt to generate 200 sentences ranging from 2 to 4 words per sentence, 150 sentences ranging from 5 to 7 words per sentence and 100 sentences ranging from 8 to 10 words per sentence. [^6] https://chatgpt.com/ Not all sentences generated by ChatGPT were correct. We discarded some sentences that were flagged by the annotators. Appendix D shows the prompts and the percentage of discarded sentences for each prompt.
Collection of Children poems
Example of the included poems: My language sings (\<لغتي تغني>), Poetry and news (\<أشعار وأخبار>), and The cat and the Eid’s hat (\<القطة وقبعة العيد>) Al-Safadi (2005); Taha-Thomure (2007).
UN
Subtitles
The Suspended Odes (Odes)
The first and the last ten verses of the ten most celebrated poems from Pre-Islamic Arabia (\<المعلقات> Mu’allaqat). All texts were extracted from Wikipedia. [^8] https://ar.wikipedia.org/wiki/\<المعلقات¿
Quran
The first Surah, the last 14 Surahs, the first 106 verses from the second Surah and the first 108 verses from the third Surah from the Holy Quran. We selected the text from the Quran Corpus Project Dukes et al. (2013). [^9] https://corpus.quran.com/
Hadith
The first 75 Hadiths from Sahih Bukhari al Bukhari (846). We selected the text from the LK Hadith Corpus [^10] https://github.com/ShathaTm/LK-Hadith-Corpus Altammami et al. (2019).
@@
Old Testament
The first 225 words from each of the first 20 chapters of the Book of Genesis Smith and Van Dyck (1865). [^13] https://www.arabicbible.com/
New Testament
[…]
@@
WikiNews
Arabic Learner Corpus (ALC)
20 L2 articles from the Arabic Learner Corpus (Alfaifi, 2015)
ArabicMMLU
6093 question and answer pairs from the ArabicMMLU benchmark dataset Koto et al. (2024).
Basic Travel Expressions Corpus (BTEC)
20 documents from the MSA translation of the Basic Traveling Expression Corpus Eck and Hori (2005); Takezawa et al. (2007).
Constitutions
The first 2000 words of the Arabic constitutions from 17 Arabic speaking countries, collected from MCWC dataset El-Haj and Ezzini (2024).
Green Library
Kalima
The first 500 words of 145 books from Kalima project. [^16] https://alc.ae/publications/kalima/
@@ 20 L2 articles from the Arabic Learner Corpus (Alfaifi, 2015)
10 manually typed editions of Majed magazine for children from 1983 to 2019.171717https://archive.org/details/majid_magazine
Qatar Arabic Language Bank (QALB)
200 online comments from the Qatar Arabic Language Bank (QALB) Corpus Mohit et al. (2014); Zaghouani et al. (2014).
@@
The opening songs of 56 animated children series from Spacetoon channel.
Zayed Arabic-English Bilingual Undergraduate Corpus (ZAEBUC)
100 student-written articles from the Zayed University Arabic-English Bilingual Undergraduate Corpus Habash and Palfreyman (2022). Some datasets are chosen because they already have annotations available for other tasks. For example, dependency treebank annotations exist for Odes, Quran, Hadith, 1001, Hayy, OT, NT, Sara,WikiNews, ALC, BTEC, QALB, and ZAEBUC Habash et al. (2022a).
@@ C.3 Readership Groups
Appendix D ChatGPT Prompts
Appendix E Additional Results
E.1 All Loss Functions
@@
E.2 Impact of Training Granularity on Readability Level Prediction
To analyze the effect of training granularity on readability level prediction, we compare two approaches: (1) training on all 19 levels and then mapping predictions to lower levels (7, 5, or 3), and (2) training directly on the target granularity (7, 5, or 3 levels). Table 11 presents the results of this comparison. Overall, the results show that training on 19 levels and mapping slightly outperforms the direct training on the target granularity.
E.3 Ensembles & Oracles
@@
Appendix F Corpus Statistics
F.1 Summary statistics of the Barec Corpus in documents
F.2 Summary statistics of the Barec Corpus in words

### v2

@@ 6.4 Ensemble Results
To further improve performance, we experiment with ensemble methods. We define the Average ensemble, where the final prediction is the rounded average of the levels predicted by the six models, and the Most Common ensemble, where the final prediction is the predicted levels’ mode. The results show that the Average ensemble performs better in terms of Distance, indicating that it tends to stay closer to the correct label. However, it struggles with exact accuracy (Acc), as averaging can blur distinctions between classes. On the other hand, the Most Common ensemble achieves higher Acc but can sometimes be misled by an incorrect majority, leading to greater deviation from the correct label.
@@
We also report an Oracle Combination, where we assume access to the best possible prediction from the six models for each sample. This serves as an upper bound on model performance. The Oracle results are significantly higher than those of individual models and are comparable to human annotators’ agreement with the unified labels (see section 6.1). This suggests that while individual models are still far from human-level performance, ensembling has the potential to push results closer to human agreement. More oracle combinations are provided in Appendix D.4. We also include more results on the impact of training granularity on readability level prediction in Appendix D.3 Finally, table 8 shows the results on the test set. We note that the trends observed in the development set persist in the test set, further validating our findings.
6.5 Error Analysis
To assess the errors in our best-performing model, we analyzed error patterns in the inter-annotator portion of the development (DEV) set. Each sentence in this subset had five human annotations, which we compared to the model’s prediction. We grouped sentences by the level of annotator agreement, from full agreement (5 out of 5 annotators) down to minimal agreement (1 out of 5). Full 5-way agreement accounts for 25% of the data. With each reduction in agreement – to 4, 3, 2, and finally 1 annotator – the cumulative coverage increases to 50%, 61%, 72%, and 87%, respectively. In other words, in 87% of the cases, the model prediction can be meaningfully compared to at least some level of human consensus. The remaining 13% fall outside this range. In 1% of these, the model’s prediction was within the span of human annotations but did not exactly match any of them. In 3%, the prediction was above the maximum annotation, and in 9%, it was below the minimum. We manually reviewed these out-of-range cases and found that the annotators were generally correct. We speculate that the model’s errors arise from limited training data, lack of contextual understanding, or insufficient modeling of linguistic features. For example, the obscure word \<عصامة> $\varsigma$$\hbar$ ‘tightly wound head dress’ may be misinterpreted as the feminine form of the proper name \<عصام> $\varsigma$SAm ‘Esam’, much like connecting \<كريم> krym ‘Kareem’ with \<كريمة> $\hbar$ ‘Kareema’. However, \<عصامة> $\varsigma$$\hbar$ is not a plausible proper name. This remains speculative, as our model is not inherently interpretable.
@@ 7 Conclusions and Future Work
This paper presented the Balanced Arabic Readability Evaluation Corpus (BAREC), a large-scale, finely annotated dataset for assessing Arabic text readability across 19 levels. With over 69K sentences and 1 million words, it is the largest Arabic corpus for readability assessment, covering diverse genres, topics, and audiences, to our knowledge. High inter-annotator agreement ensures reliable annotations. Through benchmarking various readability assessment techniques, we highlighted both the challenges and opportunities in Arabic readability modeling, demonstrating promising performance across different methods. Looking ahead, we plan to expand the corpus, enhancing its size and diversity to cover additional genres and topics. We also aim to add annotations related to vocabulary leveling and syntactic treebanks to study less-explored genres in syntax. Future work will include analyzing readability differences across genres and topics. Additionally, the tools we have developed will be integrated into a system to help children’s story writers target specific reading levels. The Barec dataset, its annotation guidelines, and benchmark results, will be made publicly available to support future research and educational applications in Arabic readability assessment.
Acknowledgments
The Barec project is supported by the Abu Dhabi Arabic Language Centre (ALC) / Department of Culture and Tourism, UAE. We acknowledge the support of the High Performance Computing Center at New York University Abu Dhabi. We are deeply grateful to our outstanding annotation team: Mirvat Dawi, Reem Faraj, Rita Raad, Sawsan Tannir, and Adel Wizani, Samar Zeino, and Zeina Zeino. Special thanks go to Abdallah Abushmaes, Karin Aghadjanian, and Omar Al Ayyoubi of the ALC for their continued support. We would also like to thank the Zayed University ZAI Arabic Language Research Center team, in particular Hamda Al-Hadhrami, Maha Fatha, and Metha Talhak, for their valuable contributions to typing materials for the project. We also acknowledge Ali Gomaa and his team for their additional support in this area. Finally, we thank our colleagues at the New York University Abu Dhabi Computational Approaches to Modeling Language (CAMeL) Lab, Muhammed Abu Odeh, Bashar Alhafni, Ossama Obeid, and Mostafa Saeed, as well as Nour Rabih (Mohamed bin Zayed University of Artificial Intelligence) for their helpful conversations and feedback.
@@ C.1 Resources
We present the corpus sources in groups of their general intended purpose.
C.1.1 Education
@@
ArabicMMLU
6,205 question and answer pairs from the ArabicMMLU benchmark dataset Koto et al. (2024).
Zayed Arabic-English Bilingual Undergraduate Corpus (ZAEBUC)
100 student-written articles from the Zayed University Arabic-English Bilingual Undergraduate Corpus Habash and Palfreyman (2022).
Arabic Learner Corpus (ALC)
16 L2 articles from the Arabic Learner Corpus (Alfaifi, 2015).
Basic Travel Expressions Corpus (BTEC)
20 documents from the MSA translation of the Basic Traveling Expression Corpus Eck and Hori (2005); Takezawa et al. (2007); Bouamor et al. (2018).
Collection of Children poems
Example of the included poems: My language sings (\<لغتي تغني>), and Poetry and news (\<أشعار وأخبار>) Al-Safadi (2005); Taha-Thomure (2007).
ChatGPT
To add more children’s materials, we ask Chatgpt to generate 200 sentences ranging from 2 to 4 words per sentence, 150 sentences ranging from 5 to 7 words per sentence and 100 sentences ranging from 8 to 10 words per sentence. [^5] https://chatgpt.com/ Not all sentences generated by ChatGPT were correct. We discarded some sentences that were flagged by the annotators. Table 9 shows the prompts and the percentage of discarded sentences for each prompt.
C.1.2 Literature
@@
A subset of 264 books extracted from the Hindawi Foundation website across different different genres. [^6] https://www.hindawi.org/books/categories/
Kalima
The first 500 words of 62 books from Kalima project. [^7] https://alc.ae/publications/kalima/
Green Library
@@
The Suspended Odes (Odes)
The ten most celebrated poems from Pre-Islamic Arabia (\<المعلقات> Mu’allaqat). All texts were extracted from Wikipedia. [^12] https://ar.wikipedia.org/wiki/\<المعلقات¿
C.1.3 Media
@@
10 manually typed editions of Majed magazine for children from 1983 to 2019.131313https://archive.org/details/majid_magazine
@@
The opening songs of 53 animated children series from Spacetoon channel.
Subtitles
WikiNews
C.1.4 References
Wikipedia
A subset of 168 Arabic wikipedia articles covering Culture, Figures, Geography, History, Mathematics, Sciences, Society, Philosophy, Religions and Technologies. [^14] https://ar.wikipedia.org/
Constitutions
The first 2000 words of the Arabic constitutions from 16 Arabic speaking countries, collected from MCWC dataset El-Haj and Ezzini (2024).
UN
C.1.5 Religion
Old Testament
The first 20 chapters of the Book of Genesis Smith and Van Dyck (1865). [^16] https://www.arabicbible.com/
New Testament
[…]
Quran
The first three Surahs and the last 14 Surahs from the Holy Quran. We selected the text from the Quran Corpus Project Dukes et al. (2013). [^17] https://corpus.quran.com/
Hadith
The first 75 Hadiths from Sahih Bukhari al Bukhari (846). We selected the text from the LK Hadith Corpus [^18] https://github.com/ShathaTm/LK-Hadith-Corpus Altammami et al. (2019). Some datasets are chosen because they already have annotations available for other tasks. For example, dependency treebank annotations exist for Odes, Quran, Hadith, 1001, Hayy, OT, NT, Sara,WikiNews, ALC, BTEC, and ZAEBUC Habash et al. (2022a).
@@ C.3 Readership Groups
Appendix D Additional Results
D.1 Confusion Matrix
Figure 3 shows the confusion matrix for the best-performing model from Stage 1: the AraBERTv2 model trained on D3Tok sentences with Cross-Entropy (CE) loss. The matrix uses F-scores to account for the unbalanced distribution of readability levels. The strong diagonal indicates a high rate of exact matches between predicted and gold labels. However, the model exhibits more disagreement at the higher, more difficult levels— likely due to the scarcity of training examples in those levels. Additionally, the model shows a tendency to under-estimate readability levels, favoring lower labels. This aligns with the patterns observed in the error analysis discussed in Section 6.5.
Figure 3: Confusion matrix of F-score across the different readability levels for the best model from stage 1.
D.2 All Loss Functions
Input Model Acc19 $\pm$1 Acc19 Dist QWK
Word SVM 36.2% 47.9% 2.03 53.4%
D3Tok SVM 37.2% 49.3% 1.92 56.5%
Word DecisionTree 27.2% 41.2% 2.50 44.2%
D3Tok DecisionTree 29.9% 44.2% 2.33 52.5%
D3Tok AraBERTv2
+CE 56.6% 69.9% 1.14 80.0%
+EMD 55.3% 70.3% 1.11 81.2%
+OLL2 35.2% 70.3% 1.25 82.0%
+OLL15 47.3% 71.1% 1.13 82.8%
+OLL1 50.8% 71.5% 1.12 81.7%
+OLL05 53.1% 68.8% 1.18 79.7%
+SOFT2 55.8% 69.8% 1.15 80.0%
+SOFT3 56.4% 69.9% 1.14 80.1%
+SOFT4 56.4% 69.9% 1.15 79.6%
+SOFT5 56.2% 69.5% 1.17 79.3%
+Reg 43.1% 73.1% 1.13 84.0%
@@ D.2 All Loss Functions
D.3 Impact of Training Granularity on Readability Level Prediction
To analyze the effect of training granularity on readability level prediction, we compare two approaches: (1) training on all 19 levels and then mapping predictions to lower levels (7, 5, or 3), and (2) training directly on the target granularity. Table 12 presents the results of this comparison. Overall, training on 19 levels and then mapping achieves slightly better performance across for 5-level and 3-level granularities compared to direct training. Moreover, the performance gap between the two approaches widens as the target granularity becomes coarser, suggesting that finer-grained supervision during training provides more informative learning signals, which translate into improved generalization when predictions are mapped into broader scales.
Train Gran Dev Gran Input Model Acc $\pm$1 Acc Dist QWK
19 7 D3Tok CE 65.9% 88.9% 0.51 79.9%
7 7 D3Tok CE 65.2% 89.5% 0.50 81.0%
19 5 D3Tok CE 70.3% 93.5% 0.37 78.3%
5 5 D3Tok CE 67.8% 93.7% 0.39 77.3%
19 3 D3Tok CE 76.5% 97.6% 0.26 74.7%
3 3 D3Tok CE 74.4% 96.9% 0.29 74.0%
@@ D.3 Impact of Training Granularity on Readability Level Prediction
D.4 Ensembles & Oracles
CE CE CE CE EMD Reg Metrics
Word Lex D3Tok D3Lex D3Tok D3Tok Acc19 $\pm$1 Acc19 Dist QWK
51.6% 65.9% 1.32 76.3%
50.1% 65.4% 1.29 77.7%
56.6% 69.9% 1.14 80.0%
53.2% 67.1% 1.24 78.6%
55.3% 70.3% 1.11 81.2%
43.1% 73.1% 1.13 84.0%
Average 46.9% 72.5% 1.11 83.4%
Most Common 56.3% 70.0% 1.13 80.4%
Oracle Combinations
62.4% 76.6% 0.88 88.4%
63.5% 76.7% 0.89 87.7%
63.2% 76.6% 0.88 88.2%
63.3% 77.9% 0.83 89.2%
62.0% 80.7% 0.77 90.8%
69.5% 82.3% 0.67 91.4%
72.0% 84.5% 0.59 92.6%
73.6% 86.6% 0.53 93.4%
75.2% 87.4% 0.50 93.8%